\begin{thebibliography}{9}
%\cite{LHCHiggsCrossSectionWorkingGroup:2016ypw}
\bibitem{LHCHiggsCrossSectionWorkingGroup:2016ypw}
D.~de Florian \textit{et al.} [LHC Higgs Cross Section Working Group],
%``Handbook of LHC Higgs Cross Sections: 4. Deciphering the Nature of the Higgs Sector,''
doi:10.23731/CYRM-2017-002
[arXiv:1610.07922 [hep-ph]].
%2034 citations counted in INSPIRE as of 18 Oct 2023

%\cite{Dawson:1990zj}
\bibitem{Dawson:1990zj}
S.~Dawson,
%``Radiative corrections to Higgs boson production,''
Nucl. Phys. B \textbf{359} (1991), 283-300
doi:10.1016/0550-3213(91)90061-2
%1005 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Graudenz:1992pv}
\bibitem{Graudenz:1992pv}
D.~Graudenz, M.~Spira and P.~M.~Zerwas,
%``QCD corrections to Higgs boson production at proton proton colliders,''
Phys. Rev. Lett. \textbf{70} (1993), 1372-1375
doi:10.1103/PhysRevLett.70.1372
%447 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Djouadi:1991tka}
\bibitem{Djouadi:1991tka}
A.~Djouadi, M.~Spira and P.~M.~Zerwas,
%``Production of Higgs bosons in proton colliders: QCD corrections,''
Phys. Lett. B \textbf{264} (1991), 440-446
doi:10.1016/0370-2693(91)90375-Z
%1059 citations counted in INSPIRE as of 01 Nov 2023

\bibitem{Spira:1995rr}
M.~Spira, A.~Djouadi, D.~Graudenz and P.~M.~Zerwas,
%``Higgs boson production at the LHC,''
Nucl. Phys. B \textbf{453} (1995), 17-82
doi:10.1016/0550-3213(95)00379-7
%[arXiv:hep-ph/9504378 [hep-ph]].
%1584 citations counted in INSPIRE as of 04 Dec 2023

%\cite{Anastasiou:2002yz}
\bibitem{Anastasiou:2002yz}
C.~Anastasiou and K.~Melnikov,
%``Higgs boson production at hadron colliders in NNLO QCD,''
Nucl. Phys. B \textbf{646} (2002), 220-256
doi:10.1016/S0550-3213(02)00837-4
[arXiv:hep-ph/0207004 [hep-ph]].
%1275 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Harlander:2002wh}
\bibitem{Harlander:2002wh}
R.~V.~Harlander and W.~B.~Kilgore,
%``Next-to-next-to-leading order Higgs production at hadron colliders,''
Phys. Rev. Lett. \textbf{88} (2002), 201801
doi:10.1103/PhysRevLett.88.201801
[arXiv:hep-ph/0201206 [hep-ph]].
%1374 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Ravindran:2003um}
\bibitem{Ravindran:2003um}
V.~Ravindran, J.~Smith and W.~L.~van Neerven,
%``NNLO corrections to the total cross-section for Higgs boson production in hadron hadron collisions,''
Nucl. Phys. B \textbf{665} (2003), 325-366
doi:10.1016/S0550-3213(03)00457-7
[arXiv:hep-ph/0302135 [hep-ph]].
%953 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Anastasiou:2014lda}
\bibitem{Anastasiou:2014lda}
C.~Anastasiou, C.~Duhr, F.~Dulat, E.~Furlan, T.~Gehrmann, F.~Herzog and B.~Mistlberger,
%``Higgs Boson GluonFfusion Production Beyond Threshold in N$^{3}LO$ QCD,''
JHEP \textbf{03} (2015), 091
doi:10.1007/JHEP03(2015)091
[arXiv:1411.3584 [hep-ph]].
%151 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Anastasiou:2014vaa}
\bibitem{Anastasiou:2014vaa}
C.~Anastasiou, C.~Duhr, F.~Dulat, E.~Furlan, T.~Gehrmann, F.~Herzog and B.~Mistlberger,
%``Higgs boson gluon\textendash{}fusion production at threshold in N$^3$LO QCD,''
Phys. Lett. B \textbf{737} (2014), 325-328
doi:10.1016/j.physletb.2014.08.067
[arXiv:1403.4616 [hep-ph]].
%198 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Anastasiou:2015vya}
\bibitem{Anastasiou:2015vya}
C.~Anastasiou, C.~Duhr, F.~Dulat, F.~Herzog and B.~Mistlberger,
%``Higgs Boson Gluon-Fusion Production in QCD at Three Loops,''
Phys. Rev. Lett. \textbf{114} (2015), 212001
doi:10.1103/PhysRevLett.114.212001
[arXiv:1503.06056 [hep-ph]].
%607 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Harlander:2009mq}
\bibitem{Harlander:2009mq}
R.~V.~Harlander and K.~J.~Ozeren,
%``Finite top mass effects for hadronic Higgs production at next-to-next-to-leading order,''
JHEP \textbf{11} (2009), 088
doi:10.1088/1126-6708/2009/11/088
[arXiv:0909.3420 [hep-ph]].
%280 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Pak:2009dg}
\bibitem{Pak:2009dg}
A.~Pak, M.~Rogal and M.~Steinhauser,
%``Finite top quark mass effects in NNLO Higgs boson production at LHC,''
JHEP \textbf{02} (2010), 025
doi:10.1007/JHEP02(2010)025
[arXiv:0911.4662 [hep-ph]].
%292 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Harlander:2009bw}
\bibitem{Harlander:2009bw}
R.~V.~Harlander and K.~J.~Ozeren,
%``Top mass effects in Higgs production at next-to-next-to-leading order QCD: Virtual corrections,''
Phys. Lett. B \textbf{679} (2009), 467-472
doi:10.1016/j.physletb.2009.08.012
[arXiv:0907.2997 [hep-ph]].
%192 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Harlander:2009my}
\bibitem{Harlander:2009my}
R.~V.~Harlander, H.~Mantler, S.~Marzani and K.~J.~Ozeren,
%``Higgs production in gluon fusion at next-to-next-to-leading order QCD for finite top mass,''
Eur. Phys. J. C \textbf{66} (2010), 359-372
doi:10.1140/epjc/s10052-010-1258-x
[arXiv:0912.2104 [hep-ph]].
%262 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Aglietti:2004nj}
\bibitem{Aglietti:2004nj}
U.~Aglietti, R.~Bonciani, G.~Degrassi and A.~Vicini,
%``Two loop light fermion contribution to Higgs production and decays,''
Phys. Lett. B \textbf{595} (2004), 432-441
doi:10.1016/j.physletb.2004.06.063
[arXiv:hep-ph/0404071 [hep-ph]].
%552 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Actis:2008ug}
\bibitem{Actis:2008ug}
S.~Actis, G.~Passarino, C.~Sturm and S.~Uccirati,
%``NLO Electroweak Corrections to Higgs Boson Production at Hadron Colliders,''
Phys. Lett. B \textbf{670} (2008), 12-17
doi:10.1016/j.physletb.2008.10.018
[arXiv:0809.1301 [hep-ph]].
%602 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Actis:2008ts}
\bibitem{Actis:2008ts}
S.~Actis, G.~Passarino, C.~Sturm and S.~Uccirati,
%``NNLO Computational Techniques: The Cases H ---\ensuremath{>} gamma gamma and H ---\ensuremath{>} g g,''
Nucl. Phys. B \textbf{811} (2009), 182-273
doi:10.1016/j.nuclphysb.2008.11.024
[arXiv:0809.3667 [hep-ph]].
%308 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Anastasiou:2008tj}
\bibitem{Anastasiou:2008tj}
C.~Anastasiou, R.~Boughezal and F.~Petriello,
%``Mixed QCD-electroweak corrections to Higgs boson production in gluon fusion,''
JHEP \textbf{04} (2009), 003
doi:10.1088/1126-6708/2009/04/003
[arXiv:0811.3458 [hep-ph]].
%467 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Anastasiou:2016cez}
\bibitem{Anastasiou:2016cez}
C.~Anastasiou, C.~Duhr, F.~Dulat, E.~Furlan, T.~Gehrmann, F.~Herzog, A.~Lazopoulos and B.~Mistlberger,
%``High precision determination of the gluon fusion Higgs boson cross-section at the LHC,''
JHEP \textbf{05} (2016), 058
doi:10.1007/JHEP05(2016)058
[arXiv:1602.00695 [hep-ph]].
%520 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Dulat:2018rbf}
\bibitem{Dulat:2018rbf}
F.~Dulat, A.~Lazopoulos and B.~Mistlberger,
%``iHixs 2 \textemdash{} Inclusive Higgs cross sections,''
Comput. Phys. Commun. \textbf{233} (2018), 243-260
doi:10.1016/j.cpc.2018.06.025
[arXiv:1802.00827 [hep-ph]].
%104 citations counted in INSPIRE as of 01 Nov 2023

%\cite{Cacciari:2015jma}
\bibitem{Cacciari:2015jma}
M.~Cacciari, F.~A.~Dreyer, A.~Karlberg, G.~P.~Salam and G.~Zanderighi,
%``Fully Differential Vector-Boson-Fusion Higgs Production at Next-to-Next-to-Leading Order,''
Phys. Rev. Lett. \textbf{115} (2015) no.8, 082002
[erratum: Phys. Rev. Lett. \textbf{120} (2018) no.13, 139901]
doi:10.1103/PhysRevLett.115.082002
[arXiv:1506.02660 [hep-ph]].
%278 citations counted in INSPIRE as of 13 Oct 2023

%\cite{Ciccolini:2007ec}
\bibitem{Ciccolini:2007ec}
M.~Ciccolini, A.~Denner and S.~Dittmaier,
%``Electroweak and QCD corrections to Higgs production via vector-boson fusion at the LHC,''
Phys. Rev. D \textbf{77} (2008), 013002
doi:10.1103/PhysRevD.77.013002
[arXiv:0710.4749 [hep-ph]].
%619 citations counted in INSPIRE as of 18 Oct 2023

%\cite{Ciccolini:2007jr}
\bibitem{Ciccolini:2007jr}
M.~Ciccolini, A.~Denner and S.~Dittmaier,
%``Strong and electroweak corrections to the production of Higgs + 2jets via weak interactions at the LHC,''
Phys. Rev. Lett. \textbf{99} (2007), 161803
doi:10.1103/PhysRevLett.99.161803
[arXiv:0707.0381 [hep-ph]].
%556 citations counted in INSPIRE as of 18 Oct 2023

%\cite{Denner:2014cla}
\bibitem{Denner:2014cla}
A.~Denner, S.~Dittmaier, S.~Kallweit and A.~M\"uck,
%``HAWK  2.0: A Monte Carlo program for Higgs production in vector-boson fusion and Higgs strahlung at hadron colliders,''
Comput. Phys. Commun. \textbf{195} (2015), 161-171
doi:10.1016/j.cpc.2015.04.021
[arXiv:1412.5390 [hep-ph]].
%188 citations counted in INSPIRE as of 13 Dec 2023

%\cite{Alwall:2014hca}
\bibitem{Alwall:2014hca}
J.~Alwall, R.~Frederix, S.~Frixione, V.~Hirschi, F.~Maltoni, O.~Mattelaer, H.~S.~Shao, T.~Stelzer, P.~Torrielli and M.~Zaro,
%``The automated computation of tree-level and next-to-leading order differential cross sections, and their matching to parton shower simulations,''
JHEP \textbf{07} (2014), 079
doi:10.1007/JHEP07(2014)079
[arXiv:1405.0301 [hep-ph]].
%7901 citations counted in INSPIRE as of 18 Oct 2023

%\cite{Frederix:2018nkq}
\bibitem{Frederix:2018nkq}
R.~Frederix, S.~Frixione, V.~Hirschi, D.~Pagani, H.~S.~Shao and M.~Zaro,
%``The automation of next-to-leading order electroweak calculations,''
JHEP \textbf{07} (2018), 185
[erratum: JHEP \textbf{11} (2021), 085]
doi:10.1007/JHEP11(2021)085
[arXiv:1804.10017 [hep-ph]].
%268 citations counted in INSPIRE as of 18 Oct 2023

\bibitem{Beenakker:2001rj}
W.~Beenakker, S.~Dittmaier, M.~Kramer, B.~Plumper, M.~Spira and P.~M.~Zerwas,
%``Higgs radiation off top quarks at the Tevatron and the LHC,''
Phys. Rev. Lett. \textbf{87} (2001), 201805
doi:10.1103/PhysRevLett.87.201805
[arXiv:hep-ph/0107081 [hep-ph]].
%526 citations counted in INSPIRE as of 04 Dec 2023

%\cite{Reina:2001sf}
\bibitem{Reina:2001sf}
L.~Reina and S.~Dawson,
%``Next-to-leading order results for t anti-t h production at the Tevatron,''
Phys. Rev. Lett. \textbf{87} (2001), 201804
doi:10.1103/PhysRevLett.87.201804
[arXiv:hep-ph/0107101 [hep-ph]].
%250 citations counted in INSPIRE as of 12 Oct 2023

\bibitem{Beenakker:2002nc}
W.~Beenakker, S.~Dittmaier, M.~Kr\"amer, B.~Plumper, M.~Spira and P.~M.~Zerwas,
%``NLO QCD corrections to t anti-t H production in hadron collisions,''
Nucl. Phys. B \textbf{653} (2003), 151-203
doi:10.1016/S0550-3213(03)00044-0
[arXiv:hep-ph/0211352 [hep-ph]].
%629 citations counted in INSPIRE as of 04 Dec 2023

%\cite{Reina:2001bc}
\bibitem{Reina:2001bc}
L.~Reina, S.~Dawson and D.~Wackeroth,
%``QCD corrections to associated t anti-t h production at the Tevatron,''
Phys. Rev. D \textbf{65} (2002), 053017
doi:10.1103/PhysRevD.65.053017
[arXiv:hep-ph/0109066 [hep-ph]].
%138 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Dawson:2002tg}
\bibitem{Dawson:2002tg}
S.~Dawson, L.~H.~Orr, L.~Reina and D.~Wackeroth,
%``Associated top quark Higgs boson production at the LHC,''
Phys. Rev. D \textbf{67} (2003), 071503
doi:10.1103/PhysRevD.67.071503
[arXiv:hep-ph/0211438 [hep-ph]].
%384 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Frixione:2014qaa}
\bibitem{Frixione:2014qaa}
S.~Frixione, V.~Hirschi, D.~Pagani, H.~S.~Shao and M.~Zaro,
%``Weak corrections to Higgs hadroproduction in association with a top-quark pair,''
JHEP \textbf{09} (2014), 065
doi:10.1007/JHEP09(2014)065
[arXiv:1407.0823 [hep-ph]].
%159 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Zhang:2014gcy}
\bibitem{Zhang:2014gcy}
Y.~Zhang, W.~G.~Ma, R.~Y.~Zhang, C.~Chen and L.~Guo,
%``QCD NLO and EW NLO corrections to $t\bar{t}H$ production with top quark decays at hadron collider,''
Phys. Lett. B \textbf{738} (2014), 1-5
doi:10.1016/j.physletb.2014.09.022
[arXiv:1407.1110 [hep-ph]].
%208 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Frixione:2015zaa}
\bibitem{Frixione:2015zaa}
S.~Frixione, V.~Hirschi, D.~Pagani, H.~S.~Shao and M.~Zaro,
%``Electroweak and QCD corrections to top-pair hadroproduction in association with heavy bosons,''
JHEP \textbf{06} (2015), 184
doi:10.1007/JHEP06(2015)184
[arXiv:1504.03446 [hep-ph]].
%338 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Demartin:2015uha}
\bibitem{Demartin:2015uha}
F.~Demartin, F.~Maltoni, K.~Mawatari and M.~Zaro,
%``Higgs production in association with a single top quark at the LHC,''
Eur. Phys. J. C \textbf{75} (2015) no.6, 267
doi:10.1140/epjc/s10052-015-3475-9
[arXiv:1504.00611 [hep-ph]].
%176 citations counted in INSPIRE as of 13 Oct 2023

%\cite{Demartin:2016axk}
\bibitem{Demartin:2016axk}
F.~Demartin, B.~Maier, F.~Maltoni, K.~Mawatari and M.~Zaro,
%``tWH associated production at the LHC,''
Eur. Phys. J. C \textbf{77} (2017) no.1, 34
doi:10.1140/epjc/s10052-017-4601-7
[arXiv:1607.05862 [hep-ph]].
%126 citations counted in INSPIRE as of 09 Oct 2023

%\cite{Frixione:2019fxg}
\bibitem{Frixione:2019fxg}
S.~Frixione, B.~Fuks, V.~Hirschi, K.~Mawatari, H.~S.~Shao, P.~A.~Sunder and M.~Zaro,
%``Automated simulations beyond the Standard Model: supersymmetry,''
JHEP \textbf{12} (2019), 008
doi:10.1007/JHEP12(2019)008
[arXiv:1907.04898 [hep-ph]].
%27 citations counted in INSPIRE as of 09 Oct 2023

%\cite{Broggio:2015lya}
\bibitem{Broggio:2015lya}
A.~Broggio, A.~Ferroglia, B.~D.~Pecjak, A.~Signer and L.~L.~Yang,
%``Associated production of a top pair and a Higgs boson beyond NLO,''
JHEP \textbf{03} (2016), 124
doi:10.1007/JHEP03(2016)124
[arXiv:1510.01914 [hep-ph]].
%68 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Kulesza:2015vda}
\bibitem{Kulesza:2015vda}
A.~Kulesza, L.~Motyka, T.~Stebel and V.~Theeuwes,
%``Soft gluon resummation for associated $t \bar{t} H$ production at the LHC,''
JHEP \textbf{03} (2016), 065
doi:10.1007/JHEP03(2016)065
[arXiv:1509.02780 [hep-ph]].
%56 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Broggio:2016lfj}
\bibitem{Broggio:2016lfj}
A.~Broggio, A.~Ferroglia, B.~D.~Pecjak and L.~L.~Yang,
%``NNLL resummation for the associated production of a top pair and a Higgs boson at the LHC,''
JHEP \textbf{02} (2017), 126
doi:10.1007/JHEP02(2017)126
[arXiv:1611.00049 [hep-ph]].
%60 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Kulesza:2017ukk}
\bibitem{Kulesza:2017ukk}
A.~Kulesza, L.~Motyka, T.~Stebel and V.~Theeuwes,
%``Associated $t \bar{t} H$ production at the LHC: Theoretical predictions at NLO+NNLL accuracy,''
Phys. Rev. D \textbf{97} (2018) no.11, 114007
doi:10.1103/PhysRevD.97.114007
[arXiv:1704.03363 [hep-ph]].
%49 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Kulesza:2018tqz}
\bibitem{Kulesza:2018tqz}
A.~Kulesza, L.~Motyka, D.~Schwartl\"ander, T.~Stebel and V.~Theeuwes,
%``Associated production of a top quark pair with a heavy electroweak gauge boson at NLO$+$NNLL accuracy,''
Eur. Phys. J. C \textbf{79} (2019) no.3, 249
doi:10.1140/epjc/s10052-019-6746-z
[arXiv:1812.08622 [hep-ph]].
%78 citations counted in INSPIRE as of 10 Oct 2023

%\cite{Broggio:2019ewu}
\bibitem{Broggio:2019ewu}
A.~Broggio, A.~Ferroglia, R.~Frederix, D.~Pagani, B.~D.~Pecjak and I.~Tsinikos,
%``Top-quark pair hadroproduction in association with a heavy boson at NLO+NNLL including EW corrections,''
JHEP \textbf{08} (2019), 039
doi:10.1007/JHEP08(2019)039
[arXiv:1907.04343 [hep-ph]].
%94 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Kulesza:2020nfh}
\bibitem{Kulesza:2020nfh}
A.~Kulesza, L.~Motyka, D.~Schwartl\"ander, T.~Stebel and V.~Theeuwes,
%``Associated top quark pair production with a heavy boson: differential cross sections at NLO+NNLL accuracy,''
Eur. Phys. J. C \textbf{80} (2020) no.5, 428
doi:10.1140/epjc/s10052-020-7987-6
[arXiv:2001.03031 [hep-ph]].
%68 citations counted in INSPIRE as of 12 Oct 2023

%\cite{Catani:2022mfv}
\bibitem{Catani:2022mfv}
S.~Catani, S.~Devoto, M.~Grazzini, S.~Kallweit, J.~Mazzitelli and C.~Savoini,
%``Higgs Boson Production in Association with a Top-Antitop Quark Pair in Next-to-Next-to-Leading Order QCD,''
Phys. Rev. Lett. \textbf{130} (2023) no.11, 111902
doi:10.1103/PhysRevLett.130.111902
[arXiv:2210.07846 [hep-ph]].
%9 citations counted in INSPIRE as of 18 Oct 2023

%\cite{Pagani:2020mov}
\bibitem{Pagani:2020mov}
D.~Pagani, I.~Tsinikos and E.~Vryonidou,
%``NLO QCD+EW predictions for $tHj$ and $tZj$ production at the LHC,''
JHEP \textbf{08} (2020), 082
doi:10.1007/JHEP08(2020)082
[arXiv:2006.10086 [hep-ph]].
%21 citations counted in INSPIRE as of 10 Oct 2023

\bibitem{Dittmaier:2003ej}
S.~Dittmaier, M.~Kr\"amer and M.~Spira,
%``Higgs radiation off bottom quarks at the Tevatron and the CERN LHC,''
Phys. Rev. D \textbf{70} (2004), 074010
doi:10.1103/PhysRevD.70.074010
[arXiv:hep-ph/0309204 [hep-ph]].
%395 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Dawson:2003kb}
S.~Dawson, C.~B.~Jackson, L.~Reina and D.~Wackeroth,
%``Exclusive Higgs boson production with bottom quarks at hadron colliders,''
Phys. Rev. D \textbf{69} (2004), 074027
doi:10.1103/PhysRevD.69.074027
[arXiv:hep-ph/0311067 [hep-ph]].
%329 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Wiesemann:2014ioa}
M.~Wiesemann, R.~Frederix, S.~Frixione, V.~Hirschi, F.~Maltoni and P.~Torrielli,
%``Higgs production in association with bottom quarks,''
JHEP \textbf{02} (2015), 132
doi:10.1007/JHEP02(2015)132
[arXiv:1409.5301 [hep-ph]].
%157 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Dicus:1998hs}
D.~Dicus, T.~Stelzer, Z.~Sullivan and S.~Willenbrock,
%``Higgs boson production in association with bottom quarks at next-to-leading order,''
Phys. Rev. D \textbf{59} (1999), 094016
doi:10.1103/PhysRevD.59.094016
[arXiv:hep-ph/9811492 [hep-ph]].
%177 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Balazs:1998sb}
C.~Balazs, H.~J.~He and C.~P.~Yuan,
%``QCD corrections to scalar production via heavy quark fusion at hadron colliders,''
Phys. Rev. D \textbf{60} (1999), 114001
doi:10.1103/PhysRevD.60.114001
[arXiv:hep-ph/9812263 [hep-ph]].
%141 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Harlander:2003ai}
R.~V.~Harlander and W.~B.~Kilgore,
%``Higgs boson production in bottom quark fusion at next-to-next-to leading order,''
Phys. Rev. D \textbf{68} (2003), 013001
doi:10.1103/PhysRevD.68.013001
[arXiv:hep-ph/0304035 [hep-ph]].
%514 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Harlander:2011aa}
R.~Harlander, M.~Kramer and M.~Schumacher,
%``Bottom-quark associated Higgs-boson production: reconciling the four- and five-flavour scheme approach,''
[arXiv:1112.3478 [hep-ph]].
%203 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Forte:2015hba}
S.~Forte, D.~Napoletano and M.~Ubiali,
%``Higgs production in bottom-quark fusion in a matched scheme,''
Phys. Lett. B \textbf{751} (2015), 331-337
doi:10.1016/j.physletb.2015.10.051
[arXiv:1508.01529 [hep-ph]].
%83 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Forte:2016sja}
S.~Forte, D.~Napoletano and M.~Ubiali,
%``Higgs production in bottom-quark fusion: matching beyond leading order,''
Phys. Lett. B \textbf{763} (2016), 190-196
doi:10.1016/j.physletb.2016.10.040
[arXiv:1607.00389 [hep-ph]].
%70 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Bonvini:2015pxa}
M.~Bonvini, A.~S.~Papanastasiou and F.~J.~Tackmann,
%``Resummation and matching of b-quark mass effects in $ b\overline{b}H $ production,''
JHEP \textbf{11} (2015), 196
doi:10.1007/JHEP11(2015)196
[arXiv:1508.03288 [hep-ph]].
%90 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Bonvini:2016fgf}
M.~Bonvini, A.~S.~Papanastasiou and F.~J.~Tackmann,
%``Matched predictions for the $ b\overline{b}H $ cross section at the 13 TeV LHC,''
JHEP \textbf{10} (2016), 053
doi:10.1007/JHEP10(2016)053
[arXiv:1605.01733 [hep-ph]].
%83 citations counted in INSPIRE as of 04 Dec 2023

\bibitem{Duhr:2020kzd}
C.~Duhr, F.~Dulat, V.~Hirschi and B.~Mistlberger,
%``Higgs production in bottom quark fusion: matching the 4- and 5-flavour schemes to third order in the strong coupling,''
JHEP \textbf{08} (2020) no.08, 017
doi:10.1007/JHEP08(2020)017
[arXiv:2004.04752 [hep-ph]].
%43 citations counted in INSPIRE as of 07 Dec 2023

\bibitem{Deutschmann:2018avk}
N.~Deutschmann, F.~Maltoni, M.~Wiesemann and M.~Zaro,
%``Top-Yukawa contributions to bbH production at the LHC,''
JHEP \textbf{07} (2019), 054
doi:10.1007/JHEP07(2019)054
[arXiv:1808.01660 [hep-ph]].
%18 citations counted in INSPIRE as of 07 Dec 2023

\bibitem{Pagani:2020rsg}
D.~Pagani, H.~S.~Shao and M.~Zaro,
%``RIP $ Hb\overline{b} $: how other Higgs production modes conspire to kill a rare signal at the LHC,''
JHEP \textbf{11} (2020), 036
doi:10.1007/JHEP11(2020)036
[arXiv:2005.10277 [hep-ph]].
%19 citations counted in INSPIRE as of 07 Dec 2023

\bibitem{Grojean:2020ech}
C.~Grojean, A.~Paul and Z.~Qian,
%``Resurrecting $ b\overline{b}h $ with kinematic shapes,''
JHEP \textbf{04} (2021), 139
doi:10.1007/JHEP04(2021)139
[arXiv:2011.13945 [hep-ph]].
%21 citations counted in INSPIRE as of 07 Dec 2023

%\cite{Denner:2011id}
\bibitem{Denner:2011id}
A.~Denner, S.~Dittmaier, S.~Kallweit and A.~Muck,
%``Electroweak corrections to Higgs-strahlung off W/Z bosons at the Tevatron and the LHC with HAWK,''
JHEP \textbf{03} (2012), 075
doi:10.1007/JHEP03(2012)075
[arXiv:1112.5142 [hep-ph]].
%230 citations counted in INSPIRE as of 13 Dec 2023

%\cite{Brein:2012ne}
\bibitem{Brein:2012ne}
O.~Brein, R.~V.~Harlander and T.~J.~E.~Zirke,
%``vh@nnlo - Higgs Strahlung at hadron colliders,''
Comput. Phys. Commun. \textbf{184} (2013), 998-1003
doi:10.1016/j.cpc.2012.11.002
[arXiv:1210.5347 [hep-ph]].
%206 citations counted in INSPIRE as of 13 Dec 2023

%\cite{Harlander:2013mla}
\bibitem{Harlander:2013mla}
R.~V.~Harlander, S.~Liebler and T.~Zirke,
%``Higgs Strahlung at the Large Hadron Collider in the 2-Higgs-Doublet Model,''
JHEP \textbf{02} (2014), 023
doi:10.1007/JHEP02(2014)023
[arXiv:1307.8122 [hep-ph]].
%102 citations counted in INSPIRE as of 13 Dec 2023

%\cite{Brein:2003wg}
\bibitem{Brein:2003wg}
O.~Brein, A.~Djouadi and R.~Harlander,
%``NNLO QCD corrections to the Higgs-strahlung processes at hadron colliders,''
Phys. Lett. B \textbf{579} (2004), 149-156
doi:10.1016/j.physletb.2003.10.112
[arXiv:hep-ph/0307206 [hep-ph]].
%680 citations counted in INSPIRE as of 13 Dec 2023

%\cite{Altenkamp:2012sx}
\bibitem{Altenkamp:2012sx}
L.~Altenkamp, S.~Dittmaier, R.~V.~Harlander, H.~Rzehak and T.~J.~E.~Zirke,
%``Gluon-induced Higgs-strahlung at next-to-leading order QCD,''
JHEP \textbf{02} (2013), 078
doi:10.1007/JHEP02(2013)078
[arXiv:1211.5015 [hep-ph]].
%254 citations counted in INSPIRE as of 13 Dec 2023

%\cite{Harlander:2014wda}
\bibitem{Harlander:2014wda}
R.~V.~Harlander, A.~Kulesza, V.~Theeuwes and T.~Zirke,
%``Soft gluon resummation for gluon-induced Higgs Strahlung,''
JHEP \textbf{11} (2014), 082
doi:10.1007/JHEP11(2014)082
[arXiv:1410.0217 [hep-ph]].
%135 citations counted in INSPIRE as of 13 Dec 2023


\end{thebibliography}

\documentclass[11pt]{report}
\usepackage{hyperref}
\usepackage{amsmath}
\usepackage{cleveref}
\newcommand{\HRule}{\rule{\linewidth}{0.5mm}}
\renewcommand*{\thefootnote}{\fnsymbol{footnote}}

\begin{document}
\hypersetup{pageanchor=false}

\thispagestyle{empty}

	\begin{flushright}
	  {\large
            \textbf{\href{https://https://cds.cern.ch/record/2886099}{LHCHWG-2024-001}}} \\[0.5cm]	
		{\large 	\textrm{May 2, 2023}} \\[2.0cm]
	\end{flushright}

	\begin{center}

	\textsc{\Large 	\href{https://twiki.cern.ch/twiki/bin/view/LHCPhysics/LHCHWG}{LHC Higgs Working Group}\footnote{\href{https://twiki.cern.ch/twiki/bin/view/LHCPhysics/LHCHWG}{\sl https://twiki.cern.ch/twiki/bin/view/LHCPhysics/LHCHWG}}} \\[0.5cm]
	\textsc{\Large 	Public Note} \\[1.5cm]
	
	\HRule \\[0.9cm]
	\textbf{\Large Ad interim recommendations for the Higgs boson production cross sections at $\sqrt{s}=13.6$ TeV} \\[1.0cm]
	\HRule \\[1.5cm]

	\textrm{
    {\large {\bf WG1 conveners:} Alexander Karlberg$^{1}$,
     Julie Malcles$^{2}$,
     Bernhard Mistlberger$^{3}$,
     Roberto~Di~Nardo$^{4}$ \\
     {\bf $\boldsymbol{ggF}$ conveners:}  Syed Haider Abidi$^{5}$, Robin Hayes$^{6}$, Alexander Huss$^1$, Stephen Jones$^{7}$,\\ 
     {\bf VBF conveners:} Gaetano Barone$^{8}$, Jiayi Chen$^{9}$, Stephane~Cooperstein$^{10}$, Silvia Ferrario Ravasio$^{1}$, Mathieu Pellen$^{11}$ \\
     {\bf $\boldsymbol{VH}$ conveners:} Hannah Arnold$^{6}$, Alessandro Calandri$^{12}$, Suman Chatterjee$^{13}$, Giancarlo~Ferrera$^{14}$, Ciaran Williams$^{15}$\\ 
     {\bf $\boldsymbol{t\bar{t}H/tH}$ conveners:} Malgorzata Worek$^{16}$, Marco Zaro$^{14}$\\ 
     {\bf $\boldsymbol{b\bar{b}H/bH}$ conveners:} Chayanit Asawatangtrakuldee$^{17}$, Tim Barklow$^{3}$, Michael Spira$^{18}$, Marius~Wiesemann$^{19}$
    }} \\[0.3cm]	
    \textit{$^{1}$CERN, Theoretical Physics Department, CH-1211 Geneva 23, Switzerland } \\
    \textit{$^{2}$IRFU, CEA, Universit\'e Paris-Saclay, Gif-sur-Yvette, F-91191, France} \\
    \textit{$^{3}$SLAC National Accelerator Laboratory, Stanford University, Stanford, CA 94039, USA } \\
    \textit{$^{4}$Universit\'{a} degli Studi and INFN Roma Tre, Via della Vasca Navale 84, I-00146, Rome, Italy}\\
    \textit{$^{5}$Brookhaven National Laboratory, PO Box 5000, Upton, NY 11973, USA}\\   
    \textit{$^{6}$Nikhef, Science Park 105, 1098 XG Amsterdam, The Netherlands }\\
    \textit{$^{7}$Institute for Particle Physics Phenomenology, Durham University, Durham DH1 3LE, UK}\\
    \textit{$^{8}$Brown University, Providence, RI 02912, USA} \\
    \textit{$^{9}$Department of Physics, Simon Fraser University, Burnaby, BC V5H0G9, Canada}\\
    \textit{$^{10}$University of California San Diego, 500 Gilman Dr, La Jolla, CA 92093, USA}\\
    \textit{$^{11}$Physikalisches Institut, Albert-Ludwigs-Universit\"at Freiburg, D-79104 Freiburg, Germany} \\
    \textit{$^{12}$ETH Zurich—Institute for Particle Physics and Astrophysics (IPA), Zurich, Switzerland}\\
    \textit{$^{13}$Institute of High Energy Physics (HEPHY), Austrian Academy of Sciences (\"{O}AW), Nikolsdorfer Gasse 18, 1050 Vienna, Austria}\\
    \textit{$^{14}$Universit\'{a} degli Studi di Milano $\&$ INFN, Sezione di Milano, Via Celoria 16, 20133 Milano, Italy}\\
    \textit{$^{15}$Department of Physics, University at Buffalo, The State University of New York, Buffalo, 14260, USA}\\
    \textit{$^{16}$Institute for Theoretical Particle Physics and Cosmology, RWTH Aachen University, \\D-52056 Aachen, Germany}\\
    \textit{$^{17}$Department of Physics, Faculty of Science, Chulalongkorn University, 10330 Bangkok, Thailand}\\
    \textit{$^{18}$Paul Scherrer Institut, CH–5232 Villigen PSI, Switzerland}\\
    \textit{$^{19}$Max-Planck-Institut f\"ur Physik, F\"ohringer Ring 6, 80805 M\"unchen, Germany}\\

\textbf{}
	\end{center}

        \thispagestyle{empty}



\mbox{}\vspace*{3em}
\begin{center}
    \textbf{Abstract}
 This note documents predictions for the inclusive production cross sections of the Standard Model Higgs boson at the Large Hadron Collider at a centre of mass energy of $13.6$ TeV. The predictions here are based on simple extrapolations of previously documented predictions published in the CERN Yellow Report "Deciphering the Nature of the Higgs Sector". The predictions documented in this note should serve as a reference while a more complete and update-to-date derivation of cross section predictions is in progress.
\end{center}

\newpage
\hypersetup{pageanchor=true}
\renewcommand*{\thefootnote}{\arabic{footnote}}
\setcounter{footnote}{0}
\setcounter{page}{1}
\chapter*{Higgs boson production cross sections at $\sqrt{s}=13.6$ TeV}
This documents collects the ad interim recommendation for the Higgs boson inclusive production cross section at $\sqrt{s}=13.6 $ TeV.
These are obtained from extrapolations based on the cross-section values computed at different value of $m_H$ and centre of mass energy in the CERN Yellow Report "Deciphering the Nature of the Higgs Sector" (YR4) (CERN-2017-002)~\cite{LHCHiggsCrossSectionWorkingGroup:2016ypw}.  
The values have been computed within the relevant subgroups of the WG1 LHC Higgs Working Group. 
Consequently, the derived numerical predictions are a direct product of the work of YR4 and the references cited therein, and it is therefore recommended both this note and the YR4 are cited whenever the interim predictions are used in experimental analyses.

\section*{ggF}
The ggF cross sections and uncertainties at different values of $m_H$ are reported in Table ~\ref{tab:ggF} and obtained with a linear interpolation of the YR4~\cite{LHCHiggsCrossSectionWorkingGroup:2016ypw} cross sections available at $\sqrt{s}=13\,\mathrm{TeV}$ and $\sqrt{s}=14\,\mathrm{TeV}$. 
The interpolation formula utilised is,
\begin{equation}
    \sigma(13.6\,\mathrm{TeV})=0.4\times\sigma(13\,\mathrm{TeV})+0.6 \times\sigma(14\,\mathrm{TeV}),
\end{equation}
where $\sigma(E)$ is the production cross section at centre of mass energy $E$. We have checked explicitly by computing a few cross sections exactly, that the error introduced by the linear interpolation is below the permille level.

The YR4 numbers incorporate results from:
\begin{itemize}
    \item QCD corrections in the heavy-top limit at NLO~\cite{Dawson:1990zj,Graudenz:1992pv,Djouadi:1991tka}, NNLO~\cite{Anastasiou:2002yz,Harlander:2002wh,Ravindran:2003um} and N$^3$LO~\cite{Anastasiou:2014lda,Anastasiou:2014vaa,Anastasiou:2015vya}.
    \item Exact finite quark mass effects at NLO~\cite{Graudenz:1992pv,Spira:1995rr}.
    \item Approximate top quark mass effects obtained in a $1/m_t$ expansion at NNLO~\cite{Harlander:2009mq,Pak:2009dg,Harlander:2009bw,Harlander:2009my}.
    \item Electroweak corrections at NLO~\cite{Aglietti:2004nj,Actis:2008ug,Actis:2008ts}
    \item Approximate mixed QCD-electroweak corrections obtained in the $m_t, m_W, m_Z \gg m_H$ limit~\cite{Anastasiou:2008tj}.
\end{itemize}
The combination procedure and error prescription are described in Ref.~\cite{Anastasiou:2016cez}, and numbers are produced using {\sc iHixs}~\cite{Dulat:2018rbf}.

\section*{VBF}

In Ref.~\cite{LHCHiggsCrossSectionWorkingGroup:2016ypw}, the state-of-the-art predictions were obtained at NNLO QCD + NLO EW accuracy.
The NNLO QCD corrections were obtained from {\sc proVBFH}~\cite{Cacciari:2015jma} while the EW corrections have been computed with the help of {\sc Hawk}~\cite{Ciccolini:2007ec,Ciccolini:2007jr,Denner:2014cla}.
Both types of corrections are obtained in the VBF approximation and do not include $s$-channel contributions.
They are combined using the following formula:
\begin{align}
\sigma^{\rm VBF} = \sigma^{\rm DIS}_{\rm NNLO\,QCD} (1+\delta_{\rm EW}) + \sigma_\gamma,
\end{align}
where $\delta_{\rm EW}$ is the relative EW correction and $\sigma_\gamma$ the photon-induced contributions.
The VBF cross sections at $13.6\,{\rm TeV}$ for different values of $m_H$ are reported in Table ~\ref{tab:VBF}.
These are obtained upon applying a cubic spline interpolation based on numbers computed in Ref.~\cite{LHCHiggsCrossSectionWorkingGroup:2016ypw} at various centre of mass energies.

\section*{VH}
The WH and ZH cross sections at different values of $m_H$ are reported in Table ~\ref{tab:WH} and  Table ~\ref{tab:ZH} respectively. These are obtained by first performing a quadratic fit to the available YR4 cross-sections at various centre of mass energies and then extracting the value for $\sqrt{s}=13.6$ TeV. The NNLO-QCD corrections were obtained with {\tt VH@NNLO}~\cite{Brein:2003wg,Altenkamp:2012sx,Brein:2012ne,Harlander:2013mla,Harlander:2014wda} and the NLO-EW with {\sc Hawk}~\cite{Denner:2011id,Denner:2014cla}.

\section*{$\boldsymbol{t\bar{t}H}$ and $\boldsymbol{tH}$}
The $t\bar{t}H$ and $tH$ ($t$-channel, $s$-channel, $tWH$) cross sections at different values of $m_H$ are reported in Table \crefrange{tab:ttH}{tab:tWH} and are obtained with a linear interpolation of the YR4 cross sections. In YR4, cross-sections
for $\boldsymbol{t\bar{t}H}$ and $\boldsymbol{tH}$ have been computed with {\sc MadGraph5\_aMC@NLO}~\cite{Alwall:2014hca,Frederix:2018nkq}, as follows:
\begin{itemize}
\item $t\bar{t}H$ production is computed at NLO QCD+EW accuracy
\cite{Beenakker:2001rj,Reina:2001sf,Beenakker:2002nc,Reina:2001bc,Dawson:2002tg,Frixione:2014qaa,Zhang:2014gcy,Frixione:2015zaa,Frederix:2018nkq}.
\item $t$-channel $tH$ is computed at NLO QCD, as documented in Ref.~\cite{Demartin:2015uha}. In particular, the central value of the cross
section is computed in a five-flavour scheme, but the quoted scale uncertainty
are the envelope of the scale uncertainty obtained both in the five- and four-flavour
scheme; hence, they account also for the flavor-scheme ambiguity.
\item $t$-channel $tH$ is computed at NLO QCD, as documented in Ref.~\cite{Demartin:2015uha}.
\item $tWH$ is computed at NLO QCD, following Ref.~\cite{Demartin:2016axk}. In particular, the ``diagram removal with interference'' (DR2)
approach, as implemented in the {\sc MadSTR} plugin~\cite{Frixione:2019fxg}, is employed to subtract contributions from resonant top quarks which appear at NLO.
\end{itemize}

It is worth to notice that, since YR4, important theoretical progress has been achieved for these processes, in particular: soft-gluon resummation 
has been performed for $ttH$ at NNLL accuracy~\cite{Broggio:2015lya,Kulesza:2015vda,Broggio:2016lfj,Kulesza:2017ukk}, 
and it has been subsequently combined with the EW corrections~\cite{ Kulesza:2018tqz,Broggio:2019ewu,Kulesza:2020nfh}. More recently, the NNLO
cross section has been computed~\cite{Catani:2022mfv}, relying on
an approximation solely for the two-loop amplitude; finally, EW corrections
have been computed for $tH$ production~\cite{Pagani:2020mov}; their effect is rather small (3\%), and furthermore they cannot be computed
for the $s$-, $t$- and $tW$ channels separately.


\section*{$\boldsymbol{b\bar{b}H}$}
The $b\bar bH$ cross sections at different values of $m_H$ are reported in Table ~\ref{tab:bbH} and obtained with a linear interpolation of the YR4 cross sections available at $\sqrt{s}=13.5$ TeV and $\sqrt{s}=14 $ TeV (For $M_H=125.4$ GeV we used the numbers for 13 TeV.). These are obtained for both the YR4 values and for the NLO+NNLLpart+$y_by_t$ predictions that are based on the following results:

\begin{itemize}
    \item QCD corrections within the 4-flavour scheme (4FS) at NLO including the interference between top- and bottom-Yukawa induced contributions~\cite{Dittmaier:2003ej,Dawson:2003kb,Wiesemann:2014ioa}.
    \item QCD corrections within the 5-flavour scheme at NLO \cite{Balazs:1998sb,Dicus:1998hs} and NNLO~\cite{Harlander:2003ai}.
    \item Santander matching of the 4FS and 5FS \cite{Harlander:2011aa}.
    \item Fully matched results of the 4FS and 5FS using the FONLL \cite{Forte:2015hba,Forte:2016sja} and NLO+NNLLpart+$y_by_t$ \cite{Bonvini:2015pxa,Bonvini:2016fgf} methods.
\end{itemize}

Since YR4 there have been additional work on the determination of the N$^3$LO corrections to bottom-induced Higgs-boson production in the 5FS that has been matched to the 4FS consistently \cite{Duhr:2020kzd}. Another study pointed out the relevance or dominance of non bottom-Yukawa induced processes as $gg\to Hg^*\to Hb\bar b$ \cite{Deutschmann:2018avk} and several other competing processes that overwhelm the SM bottom-Yukawa induced contribution \cite{Pagani:2020rsg}. The disentanglement of the bottom-Yukawa induced contribution from the other competing processes, however, might be possible by machine-learning techniques \cite{Grojean:2020ech}. In general, the bottom-Yukawa induced cross sections are relevant for BSM scenarios with an enhanced bottom Yukawa coupling as e.g.~supersymmetric extensions or a 2HDM of type II.

\section*{Acknowledgments}
This work was done on behalf of the LHCHWG.
\newpage

\begin{table}[!h]
\begin{center}
\caption{ \label{tab:ggF} ggF cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation.}
\begin{tabular}{ |l|l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{8}{|c|}{\bf ggF (N3LO QCD + NLO EW)} \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & +Th $\%$	& -Th $\%$	& TH Gaussian $\%$	& ±(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ \\
 \hline
120.00 & 5.611E+01 & +4.7 & -6.9 & ±4.0 & ±3.2 & ±1.9 & ±2.6\\ 
 120.50 & 5.571E+01 & +4.7 & -6.9 & ±4.0 & ±3.2 & ±1.9 & ±2.6\\ 
 121.00 & 5.531E+01 & +4.7 & -6.8 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 121.50 & 5.490E+01 & +4.7 & -6.8 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 122.00 & 5.451E+01 & +4.6 & -6.8 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 122.50 & 5.412E+01 & +4.6 & -6.8 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 123.00 & 5.374E+01 & +4.6 & -6.8 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 123.50 & 5.336E+01 & +4.6 & -6.8 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.00 & 5.298E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.10 & 5.290E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.20 & 5.283E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.30 & 5.275E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.40 & 5.268E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.50 & 5.260E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.60 & 5.253E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.70 & 5.245E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.80 & 5.238E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 124.90 & 5.231E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.00 & 5.223E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.09 & 5.217E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.10 & 5.216E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.20 & 5.209E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.30 & 5.202E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.38 & 5.196E+01 & +4.6 & -6.7 & ±3.9	& ±3.2 & ±1.9 & ±2.6\\
 125.40 & 5.194E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.50 & 5.187E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.60 & 5.180E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.70 & 5.172E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.80 & 5.165E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 125.90 & 5.158E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 126.00 & 5.151E+01 & +4.6 & -6.7 & ±3.9 & ±3.2 & ±1.9 & ±2.6\\ 
 126.50 & 5.115E+01 & +4.5 & -6.7 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 127.00 & 5.080E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 127.50 & 5.045E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 128.00 & 5.011E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 128.50 & 4.976E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 129.00 & 4.943E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 129.50 & 4.909E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.9 & ±2.6\\ 
 130.00 & 4.875E+01 & +4.5 & -6.6 & ±3.8 & ±3.2 & ±1.8 & ±2.6\\
 \hline
\end{tabular}
\end{center}
\end{table}

\begin{table}[!h]
\begin{center}
\caption{ \label{tab:VBF} VBF cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation.}
\begin{tabular}{ |l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{7}{|c|}{\bf VBF ( (approx.) NNLO QCD + NLO EW )} \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & +Scale  $\%$	& -Scale  $\%$	& ±(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ \\
 \hline
125	   &    4.078	&+0.5	&-0.3	&±2.1	&±2.1	&±0.5 \\
125.09 &	4.075	&+0.5	&-0.3	&±2.1	&±2.1	&±0.5 \\
125.38 &	4.067	&+0.5	&-0.3	&±2.1	&±2.1	&±0.5 \\
 \hline
\end{tabular}
\end{center}
\end{table}

\begin{table}[!h]
\begin{center}
\caption{ \label{tab:WH} WH cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation.}
\begin{tabular}{ |l|l|l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{9}{|c|}{\bf $pp\to WH$ (NNLO QCD + NLO EW ) } \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & +Scale $\%$	& -Scale $\%$	& ±(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ & $W^+H$ [pb] & $W^-H$ [pb]\\
 \hline
125	   &    1.457	&+0.4	&-0.7	&±1.8	&±1.6	&±0.9	&0.8889 	&0.5677   \\
125.09 &	1.453	&+0.4	&-0.7	&±1.8	&±1.6	&±0.9	&0.8870  	&0.5664 \\
125.38 &	1.442	&+0.4	&-0.7	&±1.8	&±1.6	&±0.9	&0.8801	    &0.5620 \\
 \hline
\end{tabular}
\vspace{1cm}
\caption{ \label{tab:ZH} ZH cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation.}
\begin{tabular}{ |l|l|l|l|l|l|l|l| }
 \hline
 \multicolumn{8}{|c|}{\bf $pp\to ZH$ (NNLO QCD + NLO EW ) } \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & +Scale $\%$	& -Scale $\%$	& ±(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ & $\sigma(gg\to ZH)$ [pb]\\
 \hline
125	   &    9.439E-01	&+3.7	&-3.2	&±1.6	&±1.3	&±0.9	&1.360E-01   \\
125.09 &	9.422E-01	&+3.8	&-3.2	&±1.6	&±1.3	&±0.9	&1.359E-01 \\
125.38 &	9.361E-01	&+3.8	&-3.2	&±1.6	&±1.3	&±0.9	&1.347E-01 \\
 \hline
\end{tabular}
\end{center}
\end{table}

\begin{table}[!h]
\begin{center}
\caption{ \label{tab:ttH} $t\bar{t}H$ cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation.}
\begin{tabular}{ |l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{7}{|c|}{\bf $\boldsymbol{t\bar{t}H}$ (NLO QCD + NLO EW) } \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & $+$Scale  $\%$	& $-$Scale  $\%$	& $\pm$(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ \\
 \hline
125	   &    $5.700 \cdot 10^{-1}$	& $+6.0$	
& $-9.3$	& $\pm 3.5$	& $\pm 3$	& $\pm 2$ \\
125.09 &	$5.688\cdot 10^{-1}$	& $+6.0$	
& $-9.3$	& $\pm 3.5$	& $\pm 3$	& $\pm 2$ \\
125.38 &	$5.638\cdot 10^{-1}$	& $+6.0$	
& $-9.3$	& $\pm 3.5$	& $\pm 3$	& $\pm 2$ \\
125.40 &	$5.634\cdot 10^{-1}$	& $+6.0$	
& $-9.3$	& $\pm 3.5$	& $\pm 3$	& $\pm 2$ \\
 \hline
\end{tabular}
\vspace{1cm}
\caption{ \label{tab:tH-t} $tH$ $t$-channel cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation. Sum of both $tH$ and $\bar{t}H$ cross sections is given.}
\begin{tabular}{ |l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{7}{|c|}{\bf $\boldsymbol{tH}$ $\boldsymbol{t}$-chan (NLO QCD) } \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & $+$Scale  $\%$	& $-$Scale  $\%$	& $\pm$(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ \\
 \hline
125	   &    $8.362 \cdot 10^{-2}$	& $+6.5$	& $-14.8$	
& $\pm 3.7$	& $\pm 3.5$	& $\pm 1.2$ \\
125.09 &	$8.353 \cdot 10^{-2}$	& $+6.5$	& $-14.8$	
& $\pm 3.7$	& $\pm 3.5$	& $\pm  1.2$ \\
125.38 &	$8.320 \cdot 10^{-2}$	& $+6.5$	& $-14.8$	
& $\pm 3.7$	& $\pm 3.5$	& $\pm 1.2$ \\
125.40 &	$8.317 \cdot 10^{-2}$	& $+6.5$	& $-14.8$	
& $\pm 3.7$	& $\pm 3.5$	& $\pm 1.2$ \\
 \hline
\end{tabular}
\vspace{1cm}
\caption{ \label{tab:tH-s}  $tH$ $s$-channel cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation. Sum of both $tH$ and $\bar{t}H$ cross sections is given.}
\begin{tabular}{ |l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{7}{|c|}{\bf $\boldsymbol{tH}$ $\boldsymbol{s}$-chan (NLO QCD)} \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & $+$Scale  $\%$	& $-$Scale  $\%$	& $\pm$(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ \\
 \hline
125	   &    $3.068 \cdot 10^{-3}$	
& $+2.4$	& $-1.7$	& $\pm 2.2$	& $\pm 2.2$	& $\pm 0.3$ \\
125.09 &	$3.064 \cdot 10^{-3}$	
& $+2.4$	& $-1.7$	& $\pm 2.2$	& $\pm 2.2$	& $\pm 0.3$ \\
125.38 &	$ 3.044 \cdot 10^{-3}$	
& $+2.4$	& $-1.7$	& $\pm 2.2$	& $\pm 2.2$	& $\pm 0.3$ \\
125.40 &	$3.040 \cdot 10^{-3}$	& $+2.4$	& $-1.7$	
& $\pm 2.2$	& $\pm 2.2$	& $\pm 0.3$ \\
 \hline
\end{tabular}
\vspace{1cm}
\caption{ \label{tab:tWH}  $tH$  $W$-associated cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation. Sum of both $tH$ and $\bar{t}H$ cross sections is given.}
\begin{tabular}{ |l|l|l|l|l|l|l|  }
 \hline
 \multicolumn{7}{|c|}{\bf $\boldsymbol{tWH}$ (NLO QCD)} \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & $+$Scale $\%$	& $-$Scale  $\%$	& $\pm$(PDF+$\alpha_{\mathrm{s}}) \%$ &	$\pm$PDF$\%$ & $\pm\alpha_{\mathrm{s}} \%$ \\
 \hline
125	       & $1.720 \cdot 10^{-2}$	& $+5.0$	& $-6.8$	
& $\pm 6.3$	& $\pm 6.1$	& $\pm 1.5$ \\
 \hline
\end{tabular}
\end{center}
\end{table}

\begin{table}[!h]
\begin{center}
\caption{ \label{tab:bbH} $b\bar bH$ cross sections at the LHC at 13.6 TeV and corresponding scale and PDF+$\alpha_\mathrm{s}$ uncertainties computed according to the PDF4LHC recommendation.}
\begin{tabular}{ |l|l|l|l|  }
 \hline
 \multicolumn{4}{|c|}{\bf $\boldsymbol{b\bar bH}$ (NNLO QCD in 5FS, NLO QCD in 4FS, Santander matching)} \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & +(Scale+PDF+$\alpha_{\mathrm{s}}$)$\%$	& $-$(Scale+PDF+$\alpha_{\mathrm{s}}$)$\%$ \\
 \hline
125.0  & 0.5269 & +20.1 & $-$24.0 \\
125.09 & 0.5257 & +20.3 & $-$24.0 \\
125.40 & 0.5213 & +20.4 & $-$24.1 \\
 \hline
\end{tabular}
\vspace*{0.5cm}

\begin{tabular}{ |l|l|l|l|l|  }
 \hline
 \multicolumn{5}{|c|}{\bf $\boldsymbol{b\bar bH}$ (NLO+NNLLpart+ybyt matching)} \\
 \hline
 $m_H$ [GeV]	& $\sigma$ [pb]  & ±(Ren.+Fact.+Matching Scales)$\%$ & +(PDF+$\alpha_s$+$m_b$)$\%$ & $-$(PDF+$\alpha_s$+$m_b$)$\%$ \\
 \hline
125	   & 0.568	& $\pm 5.0$ & +2.0 & $-$2.1 \\
125.09 & 0.566	& $\pm 5.1$ & +2.0 & $-$2.0 \\
125.40 & 0.563	& $\pm 4.9$ & +1.9 & $-$1.9 \\
\hline
\end{tabular}
\end{center}
\end{table}


\end{document}